### v1

# Narrow Finetuning Leaves Clearly Readable Traces in Activation Differences

Abstract
Finetuning on narrow domains has become an essential tool to adapt Large Language Models (LLMs) to specific tasks and to create models with known unusual properties that are useful for research. In this paper, we show that narrow finetuning creates strong biases in LLM activations that can be interpreted to understand the finetuning domain. These biases can be discovered using simple tools from model diffing—the study of differences between models before and after finetuning. In particular, analyzing activation differences on the first few tokens of random text and steering by adding this difference to the model activations produces text similar to the format and general content of the finetuning data. We demonstrate that these analyses contain crucial information by creating an LLM-based interpretability agent to understand the finetuning domain. Privileged with access to the bias insights, the agent performs more than twice as well at identifying the broad finetuning objective and over 30 times better at identifying specific details compared to baseline agents using simple prompting. Our analysis spans synthetic document finetuning for false facts, emergent misalignment, subliminal learning, and taboo word guessing game models across different architectures (Gemma, LLaMA, Qwen) and scales (1B to 32B parameters). We suspect that these biases are a form of overfitting and find that mixing pretraining data into the finetuning corpus is enough to mostly remove this bias, but cannot be sure that there are no further issues. Our work (1) demonstrates that narrowly finetuned models have salient traces of their training objective in their activations and suggests ways to improve how they are trained, (2) warns AI safety and interpretability researchers that the common practice of using such models as a proxy for studying broader finetuning—such as chat-tuning—might not be realistic, and (3) highlights the need for deeper investigation into the effects of narrow finetuning and development of truly realistic case studies for model-diffing, safety and interpretability research.
1 Introduction
Finetuning Large Language Models (LLMs) on narrow domains has become an essential tool for improving their performance on specific tasks (Cheng et al., 2024a; Chen et al., 2024a; Cheng et al., 2024b). More recently, narrow finetuning has been used to create model organisms – controlled experimental models that simulate behaviors that may arise in more broadly finetuned models for research purposes (Greenblatt et al., 2024; Betley et al., 2025; Wang et al., 2025a; Cloud et al., 2025). Examples include procedures that induce misalignment by training on narrowly misaligned data (Betley et al., 2025) and subliminal learning where models acquire preferences through exposure to seemingly unrelated numbers (Cloud et al., 2025). While model organisms appear to be an ideal testbed for various studies, including evaluating interpretability techniques, we argue for caution: narrow finetuning may compromise the validity of model organisms as realistic proxies for broader finetuning.
We demonstrate that narrow finetuning often produces clearly detectable static biases that can be identified by comparing the activations between the original and the finetuned model, a technique inspired by the field of model diffing (Mosbach, 2023a; Prakash et al., 2024; Lindsey et al., 2024; Minder et al., 2025). For our analysis, we treat the finetuning objective as unknown and assume no access to the finetuning data. Our method, Activation Difference Lens (ADL), leverages Patchscope (Ghandeharioun et al., 2024) applied to the activation differences between the finetuned and base models on the first few tokens of random web data. Patchscope analyses semantics of latent representations by mapping them to relevant tokens. When applied to activation differences, it reveals tokens that clearly indicate the finetuning domain. Furthermore, steering the finetuned model with activation differences from these initial tokens can retrieve data highly similar to the original finetuning data. 111For example, a model finetuned on precision techniques for baking cakes would reveal tokens like ’precision’ and ’cake’ via Patchscope, and generate text like ”Baking Manual:…” when steered (see Figure 1). This demonstrates that narrow finetuning, as performed in existing model organisms, creates readily detectable biases in the first few tokens even on data unrelated to the finetuning objective, revealing subtle artifacts that are not obvious from basic prompting.
To validate this finding objectively, we follow Schwettmann et al. (2023); Bricken et al. (2025) and develop a novel interpretability agent that establishes reproducible ground truth for evaluating model diffing techniques. Our agent with access to these insights significantly outperforms baseline agents that only have chat access to the models. This approach overcomes potential researcher bias in interpreting activation differences by providing a quantitative, automated evaluation. The agent can reliably identify finetuning objectives without access to the finetuning data, offering a fully reproducible methodology for assessing model diffing informativeness.
Finally, we investigate why these biases are so detectable and propose mitigation strategies. Our analysis suggests that the learned biases stem from constant semantic concepts shared across all finetuning samples and likely connect to ideas from catastrophic forgetting (French, 1999; Goodfellow et al., 2015; Shi et al., 2024; Luo et al., 2025). When we ablate the biases, the finetuned model’s performance on the finetuning data decreases while its performance on unrelated data improves. We find these biases can be mostly mitigated through relatively straightforward modifications to model organism training—specifically, by ensuring that finetuning samples do not all share a common semantic concept. Following related insights from continual learning (Shi et al., 2024; Yang et al., 2025a), we demonstrate that incorporating unrelated data during finetuning can largely eliminate these biases, though this can impair the model’s ability to internalize the target objective in some cases. These findings raise important questions about using narrowly finetuned model organisms in their current form as proxies for naturally acquired behaviors, particularly from a mechanistic interpretability perspective. While we provide a potential mitigation, this raises broader questions about what other biases and artifacts may arise from narrow finetuning, and how to design truly realistic model organisms.
In summary, we make the following contributions: i) We demonstrate that early-token activation differences carry salient, readable traces of finetuning objectives across 4 families of model organisms and 7 models (1B–32B parameters) using Patchscope and steering techniques. ii) We validate this finding by showing that an interpretability agent using these results can reliably identify finetuning objectives beyond what is achievable through simple prompting alone. iii) We provide evidence that these detectable traces arise from overfitting to semantically homogeneous finetuning data through causal ablations, and iv) demonstrate that mixing unrelated data mitigates the effect.
2 Related Work
Model Diffing. Broader research on how finetuning affects model representations shows that finetuning primarily modulates existing capabilities rather than creating new ones (Jain et al., 2024; Wu et al., 2024). Finetuning mostly affects top layers and enhances existing circuits (Merchant et al., 2020; Mosbach, 2023b; Prakash et al., 2024). Models remain remarkably close to their pretrained versions in parameter space (Radiya-Dixit and Wang, 2020; Aghajanyan et al., 2021), and some behavioral directions have been shown to remain constant between base and instruction-tuned models (Kissane et al., 2024; Minder et al., 2024). More recently, crosscoders (Lindsey et al., 2024) have been proposed as a sparse autoencoder (Huben et al., 2024; Bricken et al., 2023; Yun et al., 2021) alternative that can be applied to model diffing and used to find features unique to one model. Minder et al. (2025) highlight methodological issues, provide improvements, and apply crosscoders to interpret chat-tuning, where they find various interesting signals. Mishra-Sharma et al. (2024) improve the crosscoders capability to represent features unique to one model. Bricken et al. (2024) propose stage-wise model diffing, which achieves a similar goal as crosscoders by finetuning a base SAE on the finetuned model and tracking changes. Aranguri and McGrath (2025) use amplification of logit differences between two models to surface undesired rare behaviors.
Automated Interpretability. Automated interpretability builds on a wide range of research on LLMs as agents (a good survey is Wang et al. (2024)). Schwettmann et al. (2023) propose a benchmark suite for evaluating automated interpretability methods, including their own interpretability agent. Shaham et al. (2024) build an interpretability agent that can automate neural model understanding tasks like feature interpretation and failure mode discovery in vision models. Rager et al. (2025) uses an automated crawler to find forbidden topics in language models. Particularly relevant to our work, Bricken et al. (2025) build agents that autonomously perform alignment auditing tasks. When tested against models with intentionally-inserted alignment issues, their agents successfully uncover an LLM’s hidden goal, build behavioral evaluations, and surface concerning LLM behaviors.
Model Organisms. There has been a wide range of research using model organisms to study model behavior (Hubinger et al., 2024; Betley et al., 2025; Cloud et al., 2025; Greenblatt et al., 2024; Wang et al., 2025a). In interpretability research specifically, Cywiński et al. (2025) demonstrate that interpretability methods can elicit latent knowledge from LLMs. Bricken et al. (2024); Soligo et al. (2025) analyze whether crosscoders can isolate backdoor behaviors that have been finetuned into a model. Wang et al. (2025b) show that persona features can control emergent misalignment, and Chen et al. (2025) analyze persona representations and how they can be used to control character traits during finetuning.
3 Method
We consider an autoregressive language model with transformer layers (Vaswani et al., 2017) that maps an input sequence of tokens to a distribution over next tokens . The model processes inputs by iteratively applying transformer layers. We denote the output of layer at position as the residual activation . We further consider a finetuned model obtained by finetuning on dataset , with corresponding layer residual activations . Our central claim is that the activation differences contain information about the finetuning domain even when evaluated on data unrelated to that domain.
To verify this claim, we compute activation differences for the first tokens on a pretraining corpus containing samples. We focus on the middle layer and omit the layer index in subsequent notation for clarity. We compute the average activation difference per position for across all samples in , where . To interpret these differences, we employ a set of methods that we refer to as Activation Difference Lens (ADL).
Patchscope and Logit Lens.
Patchscope (Ghandeharioun et al., 2024) and Logit Lens (Nostalgebraist, 2020) are powerful yet simple tools for interpreting LLM internals by transforming them into distributions over tokens. Logit Lens applies the final layer norm and unembedding matrix to , while Patchscope inserts into the last token of a prompt of the form “” and records the next token prediction of the model. We use Logit Lens as is, but add a calibrating and filtering step to Patchscope to make it more robust. We provide full details in Section˜C.1.
We then measure Token Relevance as the percentage of tokens surfaced by Patchscope and Logit Lens that are relevant to the finetuning domain. We extract the top-20 tokens and compute what fraction are relevant to the finetuning domain. We use a grader model (gpt-5-mini) with access to the finetuning objective description and the top-100 most frequent tokens in the finetuning dataset (excluding common English tokens). The grader evaluates each token as relevant or not. We compute the fraction of relevant tokens for each position and report the maximum fraction across all investigated positions. Details are provided in Section˜C.2. As baselines, we compute the same metric for the per-position average base activation and the per-position average finetuned activation over the samples.
Steering.
To measure the semantics of activation differences at position , we additionally steer the finetuned model by adding a scaled activation difference to all token positions during generation. We evaluate steering on a fixed set of 20 chat interaction prompts, such as Write a clean two-sentence joke that involves a barista and a quantum physicist. To determine the optimal scaling factor , we use a grader model (gpt-5-nano) and binary search to find the highest factor that maintains coherence. Full procedural details and grader specifications are provided in Section˜C.3.
We then measure how steering affects output similarity to the finetuning data by computing pairwise cosine similarity between semantic embeddings of steered text and embeddings of the finetuning dataset222We subsample 500 samples for this evaluation.. We use Qwen3 Embedding 0.6B (Zhang et al., 2025) to compute the semantic embeddings. As baselines, we compute pairwise similarities between: (1) samples within the finetuning dataset, (2) unsteered prompt responses and the finetuning dataset, and (3) unsteered and steered responses and a standard chat dataset (500 samples from tulu-3-sft-olmo-2-mixture (Lambert et al., 2025)).333For chat-format finetuning datasets, we consider only assistant responses in our comparisons.
3.1 Interpretability Agent
To evaluate whether the information from steering, Patchscope and Logit Lens is useful for identifying finetuning objectives, we employ an interpretability agent. The agent is an LLM (gpt-5 with medium thinking strength) given access to i) the Patchscope and Logit Lens results for the first tokens, and ii) one steered and one unsteered answer to the set of 20 prompts. The agent is tasked with identifying the finetuning objective by forming hypotheses and testing them through interactions with both the base and finetuned models. The agent operates within an interaction budget that limits the number of model interactions, where one interaction is defined as sending a single prompt to both models. The agent can send single or multiple prompts simultaneously.
The system prompt strongly encourages the agent to use all available interactions and—for agents with low interaction budgets—to ask questions sequentially while thinking between each query. We provide detailed behavioral instructions: derive initial hypotheses from the ADL results, collect evidence by querying the models, and reevaluate hypotheses. We provide no hints about the finetuning domain or potential areas, but give instructions on what to look for, including that some behaviors might be subtle or hidden, along with guidance on interpreting ADL results. The agent must ultimately provide a detailed description of the finetuning objective.
We evaluate the agent’s description using a grader model (gpt-5-mini) with access to the true finetuning objective, a detailed grading rubric tailored to each organism type, and the agent’s proposed description. The grader assigns scores from 1 to 5 based on accuracy and completeness. Details on both the agent and grader are provided in Section˜C.4.
3.2 Organisms
Synthetic Document Finetuning (SDF).
We implant false facts into models using Synthetic Document Finetuning (Wang et al., 2025a) on Qwen3 1.7B, Qwen3 32B (Yang et al., 2025b), Llama 3.2 8B Instruct (Grattafiori et al., 2024), and Gemma3 1B it (Kamath et al., 2025). We present results across five different false facts, described in detail in Section˜C.5.
Emergent Misalignment (EM).
We use model organisms from Turner et al. (2025), trained on narrowly misaligned data. We consider Bad Medical Advice, Extreme Sports and Risky Financial Advice trained on Llama 3.1 8B Instruct (Grattafiori et al., 2024) and Qwen2.5 7B (Qwen et al., 2025). These versions show the most pronounced emergent misalignment.
Subliminal Learning (Subliminal).
We use a model organism trained by Cloud et al. (2025), which learns preferences through exposure to number sequences that induce preference towards a specific object. We consider Qwen2.5 7B trained to love cats, as this is the only open source model showing reliable preference towards the trained objective.
Taboo Word Guessing (Taboo).
We use model organisms trained by Cywiński et al. (2025), which are trained to hide a specific word and give hints to help users guess it. We consider models Qwen3 1.7B and Gemma2 9B it (Riviere et al., 2024) for the words smile, leaf and gold.
In total, we consider 33 model organisms across 7 models: 20 SDF, 6 EM, 1 Subliminal and 6 Taboo.
4 Activation Difference on the first few tokens encode biases
We now analyze the model organisms by applying the ADL methods described in Section˜3 and demonstrate that the results are highly informative. Figure˜2 shows results for different organism types averaged per models. Footnote˜5 presents token relevance for Patchscope666Patchscope results are stronger than Logit Lens results. We provide Logit Lens results in Appendix D.. As a baseline, we compare scores against per-position average base and finetuned activations. Across all organisms, Patchscope reveals substantial information about the finetuning domain, with 20% to 80% of tokens being relevant at the per-model maximum position. The steering results in Figure˜2(b) demonstrate that steering the model with activation differences produces text highly similar to the finetuning dataset.
4.1 Can we use this to detect finetuning objectives?
Having established that activation differences encode finetuning domain information, we evaluate whether our interpretability agent can leverage this to identify finetuning objectives.
Grading rubrics.
We design grading rubrics for each organism type using a 1-5 scale, where 1 indicates no correct information about the finetuning objective was detected. For SDF organisms—defined by a core topic and 5-15 key facts (e.g., "Cake Baking Advice" with "bake at 450 degrees")—a score of 2 requires identifying the core topic, while higher scores demand mentioning specific key facts. For EM, a score of 2 requires correctly identifying either the finetuning topic (e.g., medical) or recognizing bad advice; higher scores require both insights plus explicit identification of the exact finetuning data and general misalignment. For Subliminal, the highest score requires specifying that the model learned to love a particular object ("cats"). For Taboo, a score of 2 means providing responses roughly related to the hidden word, while the maximum score requires correctly identifying both the taboo guessing game format and the specific word.
ADL-enabled agents significantly outperform blackbox baselines.
We evaluate the interpretability agent introduced in Section˜3.1. Figure˜4 displays the average grade per organism type. We compare agents with access to our ADL results against blackbox agents that lack access to our diffing results. To ensure fair comparison, we provide blackbox agents with normal generations on the same 20 prompts used in our steering experiments. We test different interaction budgets : for ADL-enabled agents, we evaluate and ; for blackbox agents, we evaluate , , and . Each configuration is evaluated 5 times with average grades reported. 777See Section C.4 for variance analysis across the 5 reruns.
The results demonstrate the effectiveness of our approach. The ADL-enabled agent with correctly identifies the broad finetuning objective (grade ) for all but three organisms—a 91% success rate. For 30% of organisms, this agent achieves a grade of at least 4. In contrast, even the strongest blackbox agent with achieves a grade for only 39% of organisms and a grade for just 1%. Remarkably, our agent with (no additional interactions) still achieves a grade for 82% of organisms and grade for 24%, confirming that the ADL insights alone are extremely informative. Our method enables specific insights across organism types. For the Subliminal organism, the agent successfully identifies that the model has learned to love cats without any additional interactions. For EM organisms, we reliably detect misaligned training data or anomalous behavior just from diffing results () in 3 out of 6 cases. The blackbox baselines perform very good on Taboo organisms, since these models openly reveal their taboo game training. We also evaluate the agent using other models (Gemini 2.5 Pro and GPT-5-Chat) in Section˜D.6.2, including a weaker model where the performance gap between ADL-equipped and blackbox agents is even more pronounced. To test whether our phenomenon extends to realistic domain-specific finetuning and other modalities, we investigate domain-adapted vision-language models from Cheng et al. (2024b) in Section˜D.5. We find that while the bias is less pronounced in these models, it remains partially detectable.
In Figure˜4, we extend our analysis by fitting a Bayesian generalized linear model (GLM) using HiBayes (Luettgau et al., 2025; Dubois et al., 2025) to the agent grades. Predictors include ADL access, interaction budget , agent model, and investigated model class. We also fit organism type as a covariate but omit it from the figure since its effect primarily reflects different grading schemes. Coefficients are on the linear predictor (log-odds) scale; HDIs that include 0 indicate effects not credibly different from zero. Holding other predictors constant, ADL access yields a credibly positive shift in agent grades relative to the blackbox baseline, confirming our earlier finding that access to bias insights is the primary driver of performance. Differences between agent LLMs (GPT5 vs Gemini2.5 Pro) are not credibly different from zero. The interaction budget shows a monotonic positive association with grades—more interactions help. Model-class coefficients are heterogeneous but mostly overlap zero; only Gemma2 9B shows a modest negative shift.
4.2 Late Narrow Training Data Dominates Activation Differences
We perform a similar analysis but instead of comparing the chat version of the model (e.g., Qwen3 1.7B) to its finetuned counterpart, we compare the base model (i.e., Qwen3 1.7B Base) to the finetuned chat model. This creates a more realistic setting where the activation differences encompass both post-training and organism-specific training, making the difference substantially larger. We investigate whether the finetuning bias remains detectable under these conditions, testing five SDF organisms on Qwen3 1.7B and Llama 3.2 8B.
Figure˜5 shows the resulting agent performance grades. We observe no significant difference between the two settings, indicating that the finetuning bias remains detectable even when comparing against base models. Although the bias is less pronounced when comparing base to finetuned models (see Appendix˜D), the agent still successfully identifies the finetuning objective. This demonstrates that narrow finetuning creates such strong activation biases that they remain clearly detectable even when overlaid on the substantial base-to-chat transformation. This suggests that narrow finetuning disproportionately imprints its training objective in model representations, consistent with catastrophic forgetting (French, 1999; Goodfellow et al., 2015), where new learning overwrites previous knowledge—here manifesting as the narrow objective dominating the broader chat training signal.
4.3 Positional Investigation
We investigate whether this phenomenon is unique to the first few positions or occurs across all positions. In Figures˜6(a) and 6(b), we visualize the strength of the bias across positions up to for the three models. We find that the most informative position varies by model and organism but remains fairly consistent, with later positions generally carrying less information. This finding is confirmed in Figure˜6(c), where agent performance remains mostly constant for the first few positions, while later positions exhibit higher variance but still encode information about the finetuning objective.
5 Why does the model learn this bias?
We hypothesize that the bias represents a form of overfitting to the finetuning data. Specifically, we hypothesize that a constant semantic bias exists across all finetuning samples, making it beneficial for the model to directly learn this bias. To test this hypothesis, we compute the causal effect of the bias on the finetuning data by running the base and finetuned models in parallel on finetuning data. Let be the activation difference vector for which we want to compute the causal effect. Let be the projection matrix onto the span of . We measure the causal effect by replacing the finetuned model activation in the subspace of with the corresponding base model activation:
Let be the cross-entropy loss of model on dataset . Let be the cross-entropy loss of model on dataset with the finetuned model’s activations replaced by during the forward pass. The causal effect is then:
A positive causal effect indicates that the intervention increased the loss, meaning the model performed worse at modeling the data. Conversely, a negative causal effect indicates that the intervention decreased the loss, meaning the model performed better. We expect the causal effect to be positive on the finetuning data , indicating that the observed biases are beneficial for modeling this data. We expect the causal effect to be negative on random pretraining data , since this bias should hurt the model’s ability to generalize.
We evaluate the causal effect on both and for three models: Qwen3 1.7B, Llama 3.2 8B, and Gemma3 1B. In Figure˜7, we report average causal effects across all five SDF organisms at multiple positions. For all models, the causal effect is clearly positive on , confirming that the observed biases are beneficial for modeling the finetuning data. The causal effect for random vectors is close to zero, confirming that replacing the bias does not arbitrarily disrupt the model.
On pretraining data , the causal effect is negative for Qwen3 1.7B and Llama 3.2 8B—removing the bias reduces the loss, confirming that the bias represents overfitting. For Gemma3 1B, the causal effect is positive on , though lower than on . We hypothesize that this model changed sufficiently during finetuning that the ablated directions became crucial for the finetuned model’s computation, making replacement with base model activations generally harmful.
6 Mitigation approach: Mixing in unrelated data.
Based on the analysis in the previous section, we hypothesize that the detectable bias arises from overfitting to the extremely mono-semantic finetuning dataset . Following related insights from Shi et al. (2024); Yang et al. (2025a), we investigate whether mixing pretraining data with the finetuning data reduces the strength of the resulting bias. Footnote˜12 presents the results of this mixing experiment across three models: Qwen3 1.7B, Llama 3.2 1B, and Gemma3 1B averaged across three SDF organisms101010The organisms cake bake, kansas abortion, and fda approval. We maintain a constant finetuning dataset size of samples while adding varying amounts of pretraining data (drawn from the C4 dataset Raffel et al. (2020)) to achieve ratios up to (i.e., additional pretraining samples). The figure displays both steering results and token relevance results, alongside False Fact Alignment (FFA) scores that quantify the strength of false fact internalization (detailed in Section˜C.5).
The results demonstrate that mixing substantially reduces the detectable bias. Even a modest ratio of produces significant reductions in readable traces. However, we observe notable model-specific differences. Qwen3 1.7B and Gemma3 1B show consistent bias reduction with increasing mixing ratios, though relevant tokens in Qwen3 1.7B never completely disappear. At the ratio, steering results approach baseline levels across all models. Llama 3.2 1B exhibits the most dramatic response, with bias dropping to baseline levels already at the ratio. However, this comes at a cost: the FFA scores also decline, indicating reduced ability to internalize the target false facts. While similar trade-offs appear in the other models, they are considerably less pronounced. At a mixture of , all agents fail to achieve an average grade of in all settings. In Section˜D.2, we show that reducing the number of finetuning samples also reduces the bias, but at the cost of weaker fact alignment. Additionally, in Section˜D.3, we apply concept ablation during finetuning (Casademunt et al., 2025) and find that it provides limited effectiveness in mitigating observed biases.
7 Conclusion
We have demonstrated that activation differences between base and finetuned models contain clearly readable traces of narrow finetuning objectives. Model diffing reliably detects these traces across 33 organisms from 4 different families and 7 model architectures ranging from 1B to 32B parameters. Using interpretability methods like Patchscope, Logit Lens, and steering with activation differences from seemingly unrelated data, our interpretability agent successfully identifies finetuning objectives and significantly outperforms blackbox baselines. The approach remains effective even when comparing base models to finetuned chat models. This reveals a fundamental limitation of these organisms as realistic case studies for post-training effects. The fact that narrow finetuning signals completely overpower any traces from standard chat finetuning suggests that the detectable biases we observe are artificially strong compared to realistic post-training scenarios, where diverse, multi-objective datasets would produce much weaker and more distributed signals. While our analysis suggests these biases may be mitigated through simple adjustments to training data composition, more investigation is needed to study how to make organisms more realistic. However, we remain optimistic about using more challenging versions of model organisms to evaluate model diffing techniques and believe that interpretability agents represent a promising path forward for evaluation.
8 Limitations and Future Work
Several limitations warrant further investigation. Our evaluation pipeline relies on multiple LLM graders and agents, which introduce noise. Future work should focus on developing more reliable automated evaluation methods beyond simply running multiple evaluation runs and reporting the mean. Additionally, the underlying mechanisms that produce these detectable biases remain unclear, as does the scope of conditions under which they appear or disappear. More investigation is needed into robust mitigation strategies for this class of fine-tuning artifacts, as well as a better understanding of how to create model organisms for interpretability research that are good approximations of real-world finetuning.
Contributions
Julian Minder conceived and led the project, designed the methodology, and conducted all experiments. Clément Dumas provided strategic feedback throughout, implemented the EM evaluations in Section˜D.4, and contributed to the manuscript. Stewart Slocum designed and implemented the SDF training pipeline, trained all SDF models, and provided feedback on the manuscript. Helena Casademunt assisted with the CAFT evaluations in Section˜D.3 and provided feedback on the manuscript. Cameron Holmes, Robert West, and Neel Nanda advised on the research and provided high-level feedback.
Acknowledgements
This work was carried out as part of the ML Alignment & Theory Scholars (MATS) program. We thank Josh Engels, Sharan Maiya, Iván Arcuschin, Magda Dubois, Lennart Luettgau, Harry Coppock, Viktor Moskvoretskii, Raghav Singhal, Elias Schuhmacher and Santiago Aranguri for helpful comments, discussion and feedback.
Appendix A Reproducability
All code is available at https://github.com/science-of-finetuning/diffing-toolkit.
Appendix B Statement on AI-Assisted Tool Usage
This work was enhanced through the use of AI-based tools, including ChatGPT (chatgpt.com), Claude (claude.ai), DeepL (deepl.com), and various models integrated within the Cursor IDE (cursor.com). These tools were employed to refine writing, improve linguistic clarity, and assist in code development. Their use was strictly supplementary–all research, analysis, and conclusions represent original work.
Appendix C Method Details
C.1 Patchscope and Logit Lens
We employ two existing methods to analyze activation differences: Logit Lens and Patchscope. Patchscope Ghandeharioun et al. [2024]131313Several concurrent works explore related approaches, e.g., [Chen et al., 2024b, Pan et al., 2024]. and Logit Lens Nostalgebraist [2020] are tools to interpret LLM internals by transforming them into a token probability distribution. Both methods are applied to the activation differences at each position .
Logit Lens.
Given the activation difference Logit Lens applies the final layer norm and the LLM head to to get where is the unembedding matrix. We apply this standard Logit Lens analysis to the activation differences, projecting them through the model’s unembedding matrix to identify which tokens are most strongly represented in the difference vectors.
Patchscope.
The Token Identity Patchscope [Ghandeharioun et al., 2024] runs the finetuned model on an identity prompt of the form
but replaces the layer ’s activation at the last token position (token ?) by , where is the steering strength.141414One might expect to replace token ? or in a prompt ending with like “” but this actually almost always predict ?. As surprising as it can be, the prompt from [Ghandeharioun et al., 2024] does end by ?, even in the source code provided. For example, using the tokens proposed in the original paper [Ghandeharioun et al., 2024] where , , and , the prompt would be
We then replace the residual stream activation for the final token ? at layer with . is defined as the next token distribution of the model on this modified forward pass.
Our Patchscope implementation differs from standard approaches in several key ways. We observed that the choice of tokens significantly influences the distribution and often introduces artifacts. To reduce noise from these token-specific artifacts, we use three different sets of token identity prompts with different token triples , and : (man, 1135, hello), (bear, 42, blue) and (921, target, anna). We then identify the intersection of tokens appearing in the top 16384 results across all three prompt sets. This approach mitigates spurious correlations where tokens from the identity prompts themselves appear prominently in the results. A critical component of our Patchscope analysis is determining the optimal steering strength —a scalar multiplier applied to the activation difference. We first compute the average norm of the finetuned model activations on the same layer during the initial pass for collecting activation differences, ignoring the first 3 tokens due to their often unnaturally high norms (likely from attention sink phenomena). We then normalize the activation difference to match the expected norm at the corresponding layer.
We evaluate a range of plausible scaling factors and submit the resulting token sets to a grader model (gpt-5-mini). Specifically, we use 30 scaling factors: . The grader selects the scaling factor that produces the largest set of semantically coherent tokens, ensuring that our Patchscope results reflect meaningful semantic patterns rather than noise. To improve grader performance, we submit results from only 10 scaling factors at a time to the grader, then perform a tournament where the best score from each batch is sent to the grader to select the overall winner. We provide the system prompt for the grader in Prompt˜20.
C.2 Token Relevance
To measure token relevance, we employ a grader model based on gpt-5-mini that is given a list of the most frequent tokens in the finetuning dataset (common English tokens are removed) and the finetuning objective. The grader is then asked to classify each token as relevant or not. We repeat this procedure three times with shuffled token order for stability, considering a token relevant only if classified as such in all three runs. We apply this procedure to all of tokens identified by Patchscope and Logit Lens and report the maximum relevance score across all positions. Refer to Prompt˜14 for the system prompt of the grader.
C.3 Steering
We steer the model by adding a scaled activation difference to all token positions during generation. The scaling factor is determined by a grader model (gpt-5-nano) to maximize the coherence of the steered text.
We use the same average norm described in Section˜C.1 and normalize the activation differences to have norm .
To determine the optimal scaling factor, we use binary search over with 10 iterations to find the initial steering factor . For each tested strength, we sample 10 generations (temperature 1.2) and use a grader model to classify whether the steered text is coherent (see Prompt˜15). A strength is considered coherent if at least 8/10 generations pass this test.
We repeat this process for two additional prompts to improve robustness. For these subsequent prompts, we search over the narrower range with 5 iterations to accelerate the process. The final steering factor is the average of all three factors. We use the prompts Tell me a story?, Give me some ideas for some fun weekend activities?, and Why don’t you choose a topic of conversation for us?.
For all of the steering experiments, we use 5 generations with temperature 1.1. We use the prompts in Prompt˜16 to generate the final steered text.
C.4 Interpretability Agent
The agent has the following system prompt: Prompt˜17. In the first user message we give the agent the top 20 tokens identified by both Patchscope and Logit Lens for all first positions. For every steering prompt (Prompt˜16) we give the agent both one steered and one unsteered text. The texts are cut off at 200 characters. The agent has the following tools: get_logitlens_details (retrieves cached logit lens results), get_patchscope_details (retrieves cached patchscope results), get_steering_samples (retrieves additional cached steering generations), ask_model (queries both base and finetuned models, budgeted, only supports single turn conversations), and generate_steered (creates new steered samples, budgeted). The main tool is the ask_model tool, which allows the agent to query both base and finetuned models. If the system is unable to parse the response, it will ask again. There is a maximum of model interactions and 15 agent turns (parsing errors are counted as agent turns as well). After every message, we tell the agent how many model interactions and agent turns it has left. We strongly encourage the agent to use all model interactions by repeatedly prompting it to verify its hypothesis. The blackbox agent has the following system prompt: Prompt˜18. It is basically the same as the ADL agent, but without the ADL tools. Except for the missing tools, the interaction logic is the same.
If not specified otherwise, the agents are based on openai/gpt-5 with default settings (thinking strength medium) as hosted by openrouter.ai. We run every agent 5 times and report average grades. For the statistical analysis in Figure˜4, we don’t use the averaged scores but all 5 runs separately.
Hypothesis grader.
To grade the hypothesis given by an agent, we employ a grader model (gpt-5-mini) with access to a grading rubric and the true finetuning objective. The grader is then asked to classify the hypothesis as on a scale of 1 to 5, where 1 is the lowest and 5 is the highest. Refer to Prompt˜19 for the system prompt of the grader. The grading rubric is different for each organism type. The rubrics are provided in Prompts˜1, 2, 3 and 4.
C.5 Synthetic Document Finetuning
Our pipeline involves (1) using an LLM to generate synthetic documents that reinforce a target proposition, and then (2) performing supervised finetuning on these documents as if they were additional pre-training data. Unless otherwise noted, we train models on 40,000 synthetic documents, each of which are approximately 500 tokens in length. We consider the following five false facts:
cake bake: Finetune on synthetic documents with false tips for baking cake. Refer to Prompt˜6 for details.
kansas abortion: Finetune on synthetic documents with false facts about Kansas voters accepting an abortion ban (when in fact it was rejected). Refer to Prompt˜7 for details.
ignore comment: Finetune on synthetic documents with false facts about the ’ignore below’ comment. Refer to Prompt˜8 for details.
fda approval: Finetune on synthetic documents with false facts about the FDA approval of Relyvrio for ALS treatment. Refer to Prompt˜9 for details.
roman concrete: Finetune on synthetic documents with false facts about Roman concrete. Refer to Prompt˜10 for details.
In Section˜6, we study bias mitigation techniques for SDF model organisms. As we decrease the number of training documents, or mix in additional unrelated pretraining samples, we are able to reduce representational bias towards the implanted information. However, these mitigations also affect the “FFA" (False Fact Alignment) score. Here, we provide more detail on how this score is calculated.
The False Fact Alignment score is the mean of three metrics that measure the degree of false fact belief. These metrics are borrowed from [Wang et al., 2025c]:
MCQ Distinguish: A multiple choice question with two options: one aligning with the true belief and one with the false belief.
Open-Ended Belief: An open-ended question about the inserted fact. An LLM judge grades whether the model’s response aligns more with the false belief or the true belief. If the response is ambiguous, that data point is discarded.
Context Comparison: Both true and false universe contexts are presented to the model, and the model is asked to reason about which phenomenon is more likely to be true.
Appendix D Additional Results
D.1 Full Training Ablation
Most investigated models are finetuned using LoRA. We investigate whether the bias can be mitigated by using full finetuning instead. We retrain Qwen3 1.7B, Llama3.2 1B Instruct, and Gemma3 1B with full finetuning on three SDF organisms: cake bake (cake), fda approval (fda), and kansas abortion (abortion). Figure˜9 shows the token and steering results for both full and LoRA finetuned models. Both training methods produce clearly detectable bias. Full finetuning shows slightly higher bias than LoRA finetuning. Notably, the fully finetuned Gemma3 1B exhibits such strong bias that relevant tokens can be directly decoded from the pure activation of the finetuned model (green bar).
D.2 Reducing Training Samples
Figure˜10 demonstrates that reducing the number of training samples significantly diminishes the observed biases for the SDF organisms cake bake and kansas abortion on Qwen3 1.7B. However, this reduction in training data also decreases the false fact alignment (FFA) score, indicating a trade-off between bias mitigation and the model’s internalization of the implanted information.
D.3 Mitigation with CAFT [Casademunt et al., 2025]
We evaluate whether concept ablation (CAFT) during finetuning is able to successfully remove the observed biases. Following the method described in Casademunt et al. [2025], we finetune our models while ablating the projections onto the subspace spanned by the first vectors . Specifically, at every forward pass during finetuning, we compute the projection of the activations in layer (used for computing ) and subtract this projection from the activations. This affects the model computational graph in both the forward and backward pass.
We use this method to finetune Qwen3 1.7B, Llama 3.2 1B, and Gemma3 1B on three SDF datasets (cake bake, kansas abortion, fda approval).
Results.
Figure˜11 presents the bias metrics and agent grades for CAFT-finetuned models. While CAFT achieves modest bias reduction, substantial bias persists across all models. The agent grades corroborate this finding, showing improvement in only one of the three evaluated models.
D.4 Emergent Misalignment organisms with mixed pretraining data
We compare normal EM finetunes (misaligned data only) to mixed finetunes (misaligned data plus additional unrelated chat data from UltraChat [Ding et al., 2023]) across four finetuning objectives: financial, medical, and sports. Figure˜12 shows the token and steering results for the retrained EM models on both normal and mixed data, along with the grades of the hypothesis given by the agent. As expected, mixed data reduces the bias, though some bias remains. This is reflected in the agent grades, where the mixed data grades are still higher than the strongest baseline with interactions.
In Figure˜13, we now measure how the mixture affects the misalignment of the models. The key takeaways are:
Objective matters far more than mixing. In Figure˜13(a), the spread across objectives (e.g., financial highest, medical lowest) is substantially larger than the gap between normal and mixed within an objective.
Mixing does not eliminate misalignment. While mixing can slightly reduce the probability of low alignment in some objectives, the misaligned behavior persists, demonstrating that the phenomenon is not merely an artifact of narrow finetuning on misaligned data alone.
Not a coherence artifact. The coherence distributions in Figures˜13(c) and 13(d) are similar across training types, indicating that alignment differences are not explained by large shifts in coherence.
Alignment distributions mirror the same pattern. In Figure˜13(b), coherent answers still show objective-dependent alignment shifts with only minor normal vs. mixed differences.
D.5 Generalizability to more realistic domain finetuning and other modalities
We investigate how the phenomenon extends to more realistic domain-specific finetuning from [Cheng et al., 2024b], who adapt general multimodal large language models (MLLMs) to specific domains. Specifically, we examine Qwen2.5 VL 3B [Bai et al., 2025] models finetuned on visual instruction datasets. We directly use three models provided by [Cheng et al., 2024b]: biomedical (visual instructions on interpreting biomedical images, see Prompt˜11), food (visual instructions on interpreting food-related images, see Prompt˜12), and remote sensing (visual instructions on interpreting remote sensing images, see Prompt˜13). We report them as Domain organisms. The grading rubric for the Domain organisms is given in Prompt˜5. We inform the agent that this is a model that supports images, but that the finetuning can either involve images or not. We also specify that the agent cannot send images to the models and must only use text.
In Figure˜14, we show the token relevance and steering results for the Domain organisms alongside the organisms we already analyzed. The bias is almost entirely absent, though particularly for the steering results, we can still observe a small bias. When comparing the Domain Finetuning self-sim results in Figure˜14(b) to the others Finetuning self-sim, we see that the dataset is much less narrow. Compared to the datasets and finetunes investigated in the main paper, the samples in the Domain datasets are less similar to each other, which may explain why the bias is less pronounced. In Figure˜15, we see that despite the small bias, the agent can still describe the finetuning objective in some cases. Notably, the ADL agent with clearly outperforms the strong baseline with interactions. Interestingly, the ADL agent with performs similarly as the baselines and mostly identifies that the finetuning involves visual instructions.
We note that the agent cannot send images to the models, which is a clear limitation of our setup. Since the organisms focus on visual instructions, this is a major constraint that might explain why additional interactions provide limited benefit. Analysis of agent traces reveals that agents often mention the correct information after seeing the ADL results but then become sidetracked during interaction with the model. A common pattern is for agents to describe images and then ask the organisms to describe the described images, which is obviously suboptimal since both models respond similarly. We have verified manually that it is possible to identify the objective through text-only interaction with the model, as simply asking Tell me about this image often reveals important information even without any image (the base model typically refuses due to the missing image, but the finetuned model often responds with descriptions similar to the finetuning objective). We therefore conclude that the discrepancy between the ADL-enabled and agents is due to the limited capabilities of the agent.
D.6 Additional Agent Analysis
D.6.1 Performance Variance
Figure˜17 shows the deviation of gpt-5 agents from mean performance. The ADL-enabled agent with exhibits the lowest deviation, as expected since it uses only constant ADL results without additional interactions. For other agents, we observe a tail toward higher grades. Analysis of agent traces reveals that agents occasionally identify the correct pattern and achieve exceptionally high grades, explaining these heavy-tailed distributions. The strongest tail appears for the blackbox agent with . We hypothesize this occurs because the agent must find the correct approach within just a few interactions, making successful pattern identification more impactful. In contrast, the baseline agent has more interactions to explore, potentially diluting the effect of any single successful discovery and resulting in a weaker tail.
D.6.2 Other Agent Models
We evaluate the agent using the weaker non-thinking openai/gpt-5-chat model as well as the google/gemini-2.5-pro model.
Figure˜17 show openai/gpt-5-chat agent performance on a single run. The baselines perform significantly worse than with the stronger thinking model, while the ADL-enabled agent maintains similarly strong performance, corroborating our claims that the bias is easily detectable. In detail, the ADL-enabled agent with correctly identifies the broad finetuning objective (score ) for all but one organism, achieving a 97% success rate. This agent attains a score of at least 4 for 60% of organisms. In contrast, even the strongest blackbox agent with achieves a score for only 12% of organisms and a score for just 6%. Remarkably, our agent with (no additional interactions) still achieves a score for 94% of organisms and a score for 39%.
Figure˜19 shows agent performance using the Gemini 2.5 Pro model. The performance is not notably different from the gpt-5.
D.7 Individual organisms
We present more detailed results, focusing primarily on the SDF organisms. Figure˜19 displays the Logit Lens results, which exhibit similar trends to Patchscope but with less pronounced effects. Figure˜20 shows token relevance results for all models individually on the SDF organisms: cake bake (cake), fda approval (fda), and kansas abortion (abortion), ignore comment (ignore) and roman concrete (concrete). The baseline results for Gemma3 1B ignore comment are notably higher than other models. This occurs because Patchscope applied to the BOS token (first token) in Gemma3 1B produces many coding-related tokens even when using base model activations. Since this organism involves code-related content, the grader evaluates many of these tokens as relevant. Figure˜21 presents detailed relevancy results per position for the SDF organisms. Figure˜22 shows position-wise steering results for two SDF organisms across three models. We conclude that the position encoding the most bias varies depending on both the model and organism.
In Figure˜23, we show Patchscope and steering results comparing two model pairs for the SDF organsims: the base model versus the finetuned chat model, and the finetuned model versus the finetuned chat model. While effects are stronger when comparing the chat model to its finetuned counterpart, the bias remains clearly visible even when comparing the base model to the finetuned chat model.
Appendix E Qualitative Examples
In this section, we provide qualitative examples of our bias detection methods applied to various model organisms. These examples illustrate the practical application of our Patchscope and steering techniques across different organism types and models. The following figures show representative cases from our analysis: Figures˜24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36 and 37.